prove that sin (degree(m)) is an algebraic number of any integer m.



1
Expert's answer
2022-03-19T02:38:37-0400

180o"\pi" radian

mo= m"\pi"/180

em"\pi"i/180 = cos ( m"\pi"/180)+ i sin (m"\pi" /180)

(em"\pi"i/180)180 = cos (m"\pi") + i sin (m"\pi")

(-1)m =+1= -1

cos (m"\pi"/180) + i sin (m"\pi"/180)

cos (m"\pi"/180) - i sin (m"\pi"/180)


2cos (m"\pi"/180) is an algebraic integer

Hence, cos (m"\pi"/180) becomes an algebraic integer

i sin (m"\pi"/180) is an algebraic integer

sin (m"\pi"/180) also becomes an algebraic integer


"\therefore" sin (degree(m)) is an algebraic number